# Deferred Poisoning: Making the Model More Vulnerable via Hessian Singularization

### v2

@@ Abstract
Index Terms:
Poisoning Attack, Robustness, Hessian.
I Introduction
Deep learning models have achieved remarkable achievements in fields involving Computer Vision [1] to Natural Language Processing [2] even Artificial Intelligence Generated Content [3]. Their success is largely attributed to the proliferation of large-scale datasets, such as ImageNet [4], COCO [5], and CelebA [6]. Nevertheless, recent studies have highlighted the potential hazards posed by poisoning attacks which can undermine the integrity of deep learning models by introducing malicious data into training datasets [7, 8, 9].
@@ I Introduction
Most prevalent poisoning attacks [10, 11, 12, 13] share a common limitation: their malicious intentions are highly apparent. The victim may notice a significant disparity in the performance of the model between the training and validation datasets. In response, strategies such as adversarial training [14], appropriate data preprocessing methods [15], or the elimination of anomalous [16] data can be employed as defensive measures against these attacks.
In this paper, we reveal a more threatening method of poisoning attack namely Deferred Poisoning Attack (DPA). As the term "Deferred" implies, this attack does not disrupt the training process, allowing the model to maintain normal performance over the validation set. Instead, the "toxicity" of the attack manifests by undermining the model’s robustness at the deployment stage. Fig. 1 illustrates the scenario of our attack. When a recognition system (e.g. the one deployed on an autonomous driving platform) is trained using clean datasets, it functions effectively on normal targets and demonstrates relative robustness against natural noise or adversarial perturbation generated by evasion attacks such as FGSM [17], PGD [18], etc. In contrast, when the training dataset is tainted by our proposed method, the recognition system built upon the corrupted datasets operates as usual on clean inputs but exhibits high sensitivity to adversarial perturbations or even natural noise. The threat posed by DPA is evidently more severe than the typical poisoning attacks. The developers of the recognition system are unable to detect anomalies through the detection of irregularities in the training process. Once deployed in vehicles, the system would be easily misled by natural factors such as fog, rain, and varying lighting conditions. Additionally, adversaries may deceive the model at a low cost, using subtle adversarial perturbations.
We fulfill the purpose of DPA by forcing the model trained over the contaminated dataset to converge to a similar point as the one trained over the clean dataset, making the poisoned model perform normally on the validation dataset. On the other hand, we enlarge the local curvature of the poisoned model around each sample in the training dataset to amplify the sensitivity of the poisoned model. Fig. 2 illustrates the above motivation. A large local curvature (the red curve) results in a significant increase in model loss with a small perturbation of a given sample. In contrast, a small local curvature (the blue curve) enables the model loss to remain stable even with a large perturbation. Formally, a large local curvature implies that the Hessian matrix is ill-conditioned with a large conditional number [19]. Along this line, our DFA generates poisoned samples to induce the model trained on this contaminated dataset to become singularization (a large conditional number.) with respect to the input samples.
@@
  • We validate the generality of the deferred poisoning attack across a broader range of scenarios.
@@
II Related work
II-A Poisoning Attacks
Poisoning attacks aim to insert malicious samples into training datasets, thereby compromising the ability of the resulting models [20, 21, 22, 23]. Huang et al. [11] first proposed a method for calculating perturbations by minimizing empirical risk, which impedes the models’ ability to extract valuable information from images. Fu et al. [24] made a more robust improvement by minimizing adversarial training loss to obtain a more robust set of unlearnable examples. He et al. [25] proposed Sharpness-Aware Data Poisoning Attack (SAPA) to address the limited effectiveness of existing data poisoning attacks in the face of the uncertainty in the retraining process of deep neural networks, by leveraging the sharpness of the loss landscape of DNNs to optimize the poisoning effect and enhance the effectiveness and versatility of data poisoning attacks. However, previous poisoning attacks directly compromised the integrity of the models, making it easy for us to detect anomalies in the model’s performance on the validation set during the training phase, which then allows us to take defensive measures. In this paper, we identify a form of deferred poisoning that appears normal during the model training phase yet renders the poisoned model extremely vulnerable. This kind of stealthiness can mislead researchers, thereby posing a serious threat to their work.
II-B Backdoor Attacks
Backdoor attacks clandestinely insert triggers into the training data, enabling the model to function normally until the trigger is encountered, which induces an aberrant performance [26, 27, 28, 29, 30]. Turner et al. [31] first studied the potential backdoor vulnerabilities in machine learning paradigms and training processes, then proposed the BadNet attack algorithm, which makes the model learn the relationship between the trigger and a specific label. Turner et al. [32] first proposed label-consistent backdoor attacks, which disrupted the original features of the image with specific noise, which caused the model to focus on the features of the trigger rather than the original features of the image. Nguyen and Tran [33] designed a more advanced input-aware dynamic backdoor attack, which added specific triggers to each input, breaked the traditional assumption that the trigger is independent of the data. Nguyen and Tran [34] achieved an imperceptible attack by introducing subtle distortions into the image, posed a significant threat to the security of deep learning models. Traditional backdoor attacks are hard to detect but require very specific triggers to cause errors in models, which are rarely seen in real life. The DPA we introduce makes models more sensitive to minor disturbances and natural noise, fitting real-world situations better.
II-C Model Generalization
Zhao et al. [35] proposed the gradient of the loss function with respect to the input in the vicinity of a local optimum approximates the Lipschitz constant according to the Lagrange Mean Value Theorem, thereby controlling the sensitivity of the loss function to input perturbations. The smaller the gradient, the less the loss is affected by perturbations, and the stronger the model’s generalization capability. In other words, the flatter the loss curve is in the vicinity of a local optimum, the stronger the model’s generalization capability. Inspired by the relationship between the geometry of the loss landscape and generalization, Foret et al. [36] introduced the Sharpness-Aware Minimization (SAM) method to improve the generalization ability of the model by simultaneously minimizing the loss value and the loss sharpness, aiming to solve the problem of insufficient generalization ability in model training. In this paper, we introduce the Deferred Poisoning Attack (DPA), which, at first glance, seems to oppose strategies designed to improve model generalization. Typically, generalization techniques focus on the gradient or the Hessian matrix of the loss function concerning model parameters. In contrast, DPA uniquely computes the Hessian matrix of the loss function with respect to the input data. Our forthcoming experiments will assess the efficacy of generalization-boosting methods, such as SAM, in countering DPA.
III Deferred poisoning attack
In this section, we introduce the principles of deferred poisoning attack, detailing how to generate such perturbations and providing both the necessary theoretical and empirical analysis.
III-A Problem statement
@@
We assume that adversaries have access to the entire training dataset but are unaware of the target model’s parameter structure or any output, and do not interfere with the target model’s training process.
@@
[…]
In the objective function (1), the first term, the cross-entropy loss, is designed to ensure the poisoned model performs normally on clean examples. The second and third terms are dedicated to updating the poisoning perturbation $\delta$. Notably, our approach differs significantly from existing poisoning attack methods [37, 38, 39], which typically aim to increase the model loss on poisoned examples so as to degrade the model’s performance on the validation dataset. Our proposed loss still allows the model to converge on the poisoned examples but introduces a unique design: the Hessian Singularization term $Q(f_{\theta}(\hat{x}),y)=\textit{tr}(H^{T}H)$, where $H$ is the Hessian matrix of the loss function with respect to the input $\hat{x}$. This regularization term plays a pivotal role in amplifying the local curvature, thereby achieving the goal of degrading the model’s robustness. The value of $Q(f_{\theta}(\hat{x}),y)$ is positively related to the curvature around the sample $\hat{x}$. More of the relevant theoretical derivation about the Hessian Singularization term will be discussed in Session III-C. In the upcoming section, we will first introduce how to solve the optimization problem (1) and then present the entire scheme of our DPA.
III-B Generating perturbation
Empirically, to ensure that the optimization of the problem (1) converges effectively, we divide this problem into two alternating iterative sub-optimization problems, which have been widely adopted in the research area of poisoning attack [40, 12]. Specifically, as shown in Fig. 3, the first phase is the model training stage. We consider the following sub-problem
$\theta\leftarrow\mathop{\arg\min}_{\theta}\bigg{[}\mathcal{L}(f_{\theta}({x}),y)+\mathcal{L}(f_{\theta}(\hat{x}),y)\bigg{]}$ (2)
@@ III-B Generating perturbation
III-C Hessian singularization
In this section, we will explain why the Hessian Singularization could impact the robustness of deep models. Consider the loss function of a model denoted by $l$ and a given input $x$ and assume that this function is strictly convex around a small neighbor of the $x$. For any small perturbation $h$ around $x$, the loss function satisfies the following conditions [19]:
@@ III-C Hessian singularization
where $tr()$ represents the trace of a matrix (see [41] for the details of this inequality). Consequently, this leads to the proposed Hessian Singularization regularization term $Q(f_{\theta}^{*}(x+\delta),y)=\text{tr}(H^{T}H)$ in the objective function (1).
@@
III-D Relaxation of Hessian singularization term
@@
The new strategy avoids directly computing the Hessian matrix and significantly reduces the computational cost to approximately 0.008 seconds and still maintains the effectiveness of our DPA. Following is a specific analysis of computational complexity:
Time complexity
  • $\bullet$
    Stage 1. Train Model: Update model parameters: $O(MP)$ [42]
  • $\bullet$
    Stage 2. Update perturbation
    • $\bullet$
      Step 1. Perform a forward pass to obtain the Loss: $O(MP)$
    • $\bullet$
    • $\bullet$
    • $\bullet$
      Step 4. Compute $\text{tr}(H^{T}H)$ or $\|Hv\|_{F}$: $O(P)$
    • $\bullet$
      Step 5. Backpropagate and update $\epsilon$: $O(P)$
Space complexity
IV Experiments
IV-A Experiment setting
Datasets & Models
We generate perturbation on the CIFAR10 [43], SVHN [44], and TinyImageNet [4] datasets respectively. Subsequently, we use the VGG16 [45], ResNet18 [46], and ResNet50 [46] models to train on the CIFAR10 and SVHN datasets. For the TinyImageNet, we use the ResNet18, ResNet50, and DenseNet121 [47] models for training. It should be noted that the TinyImageNet is a subset of the ImageNet dataset.
Evaluation metric
The major claim of our work is that models trained over a dataset poisoned by our DPA would exhibit a lower robustness against adversarial perturbation or even natural noises. Thus we quantize the robustness of a victim model by adopting the measure defined by [48]. That is,
where $\hat{r}(x)$ is the minimal perturbation for a successful attack, and $\mathscr{D}$ denotes the validation set. $p$ is $2$ (e.g., DeepFool [48]) or $\infty$ (e.g., FGSM and PGD). A smaller $\hat{\rho}(f)$ indicates a model is more vulnerable.
On the other hand, we emphasize that our DPA could significantly decrease the degradation of the model performance on clean examples, exhibiting the stealthiness of our DPA. To this end, we apply traditional poisoning methods such as EM [49], REM [50], LSP [51], and AR [52], as well as our proposed method to contaminate datasets [15]. We assess the classification accuracy of the poisoned models, the closer the accuracy is to that of the clean models, the more effective the stealth of the poisoning attack method.
IV-B Effectiveness of deferred poisoning attack
As shown in Table I where $\hat{\rho}_{F}$, $\hat{\rho}_{P}$ and $\hat{\rho}_{D}$ represent the robustness of victim models attacked by FGSM, PGD, and DeepFool respectively, we can observe that the ACC of both the clean models and the poisoned models are almost the same. For example, the 4th and 5th rows, corresponding to the 1st to 4th columns, illustrate the performance of the model trained with ResNet18 on TinyImageNet. After being subjected to DPA, the poisoned model maintains the same accuracy as the clean model on the validation set, both at 0.72. However, the robustness of the poisoned model against FGSM, PGD, and DeepFool adversarial attacks has significantly decreased from 0.74, 0.50, and 0.28 to 0.48, 0.42, and 0.09, respectively. Similarly, by examining the data for other models and datasets in the table, we arrive at the same conclusion.
TABLE I: The comparison of the accuracy and robustness(%) of the clean model and poisoned model against different kinds of evasion attacks.
@@ IV-B Effectiveness of deferred poisoning attack
Transferability
It is important to note that, in these examples from Table I, perturbations are generated based on the VGG16 model that was pretrained on the full ImageNet dataset. However, the toxic effects of these perturbations are not only effective on the VGG16 model (white-box scenario) but also transfer to other models such as ResNet18, ResNet50, and DenseNet121 (black-box scenarios). This indicates that DPA possesses excellent transferability, thereby making it highly suitable for black-box attack scenarios. The poisoned model being simultaneously affected by adversarial samples and natural noise indicates that DPA fundamentally reduces the model’s robustness to arbitrary perturbations. This corroborates our claim that by increasing the curvature of the loss function around local optima, the model becomes more vulnerable.
Generalized scenario
Adversarial examples do not frequently manifest in real-world scenarios. According to our claim, DPA fundamentally reduces the model’s robustness to arbitrary perturbations, that’s means poisoned model not only simultaneously affected by adversarial samples but also natural noise. To substantiate this, we introduce random noise sampled from different distributions to the test images to simulate more general conditions as shown in Fig. 5.
Figure 5: Image after adding different random noise
We experimented with natural noise from various distributions:
Table II compares the classification accuracy of clean and poisoned models on CIFAR10 with different sampled random noise injections. It is evident that the poisoned models are significantly more sensitive to random noise, indicating that our method can render the models vulnerable. For example, in the second column, when perturbing VGG16 with Gaussian noise, the clean model achieves a high classification accuracy of 0.80, remaining virtually unaffected by the noise. However, when the same Gaussian noise is used to disturb the poisoned model, the classification accuracy drops to only 0.45. This indicates that DPA renders the model considerably vulnerable even to natural noise, and poses a significant potential threat to the field of artificial intelligence security.
TABLE II: The impact of natural noise on the poisoned and clean models.
Model Noise Clean Gaussian Poisson Speckle Rayleigh
VGG16 Clean 0.89 0.80 0.71 0.72 0.82
Poisoned 0.75 0.45 0.47 0.48 0.44
ResNet18 Clean 0.81 0.60 0.63 0.62 0.66
Poisoned 0.76 0.41 0.49 0.41 0.43
ResNet50 Clean 0.78 0.63 0.61 0.65 0.68
Poisoned 0.73 0.37 0.50 0.38 0.42
IV-C Stealthiness of poisoned samples
Similar to traditional poisoning attacks, we aim for the poisoned samples generated by DPA to be imperceptible to the human eye. In this section, we verify the low attack cost characteristic of DPA and thereby achieve the stealthiness of the poisoned samples.
DPA has a significant advantage over traditional data poisoning methods, it requires only a minimal perturbation cost, such as $\epsilon=3/255$, to achieve excellent results. In contrast, data poisoning methods with $l_{\infty}$ norm constraints, such as EM and REM, do not demonstrate effectiveness at this level of perturbation.
By contrasting the classification accuracy of adversarial samples between the poisoned and clean models, we present the effectiveness of DPA in a more intuitive form. Table III presents the classification accuracy of the model against different adversarial samples after being subjected to poisoning attacks with varying perturbation radii, it shows that the EM and REM methods do not fully take effect when $\epsilon$ are as low as $3/255$ or even $5/255$ across different models, at least $\epsilon=8/255$ to exhibit aggressiveness. In contrast, our method shows excellent offensiveness even at the scale of $\epsilon=3/255$. For instance, in rows 2 to 4 of Table III for the VGG16 model trained on CIFAR10, when $\epsilon=3/255$, the models trained with EM and REM maintain high classification accuracy and accuracy post-adversarial attack. However, the model poisoned by DPA exhibits extremely low accuracy post-adversarial attack, indicating that the models are very vulnerable and can be significantly impacted by adversarial examples due to the approach mentioned. This also suggests that the attack cost of DPA is much lower than that of traditional poisoning attacks, while also ensuring the stealthiness of our method.
TABLE III: The performance of various poisoning attacks at different attack scales.
@@ IV-C Stealthiness of poisoned samples
Aiming to assess stealthiness without compromising the efficacy of poisoning attacks, we set the $l_{\infty}$ norm attacks’ parameter $\epsilon$ to $8/255$ for traditional poisoning attacks, such as EM and REM. For $l_{2}$ norm attacks, such as AR and LSP, we set $\epsilon$ to $1$. However, for DPA, an $\epsilon$ setting of $3/255$ is sufficient to ensure aggressiveness.
IV-D Comparison with backdoor attacks
@@
Similar to DPA, models attacked by backdoor attacks behave normally during the training phase. Backdoor attacks are executed by introducing specific backdoors into the training samples, which prompts the model to learn the relationship between the backdoor and a designated label. This leads to the model exhibiting abnormal behavior during the testing phase when it encounters the trigger. However, DPA does not implant specific Trojans, more generally, its goal is to make poisoned models particularly vulnerable, primarily by making them more susceptible to adversarial attacks even random noise.
TABLE IV: The accuracy and robustness(%) of VGG16 after training CIFAR10 with backdoors implanted.
Methods Robustness ACC ASR $\hat{\rho}_{F}\downarrow$ $\hat{\rho}_{P}\downarrow$ $\hat{\rho}_{D}\downarrow$
@@
[…]
@@
Table IV demonstrates that even when the attack success rate (ASR) is very high, backdoor attacks do not render the model more vulnerable. This indicates that the aggressiveness of deferred poisoning attack is insidious and unique, it can make the model more vulnerable while ensuring normal model performance, bringing unforeseen potential harm to machine learning.
IV-E Robustness analysis
@@
We conduct various data augmentation techniques [53, 54] on data-infused with deferred poison. By observing the performance of models trained on these augmented datasets, we assess the robustness of DPA. Specifically, in addition to augmenting the images, we compare the effects of models both before and after they utilize weights pre-trained on ImageNet.
@@
Table V displays the performance of deferred poisoning attacks when facing various data augmentation techniques with $\epsilon$ set to $3/255$. It is evident that our method can still render models vulnerable under most data augmentation methods as well as with pre-trained models, indicating that our approach possesses strong robustness.

### v3

@@ Abstract
Codehttps://github.com/Anson-He/DPA
Introduction
Deep learning models have achieved remarkable achievements in fields involving Computer Vision (resnet) to Natural Language Processing (transformer). Their success is largely attributed to the proliferation of large-scale datasets, such as ImageNet (imagenet), COCO (coco). Nevertheless, recent studies have highlighted the potential hazards posed by poisoning attacks which can undermine the integrity of deep learning models by introducing malicious data into training datasets (better-safe-than-sorry; fowl2021adversarial; tao2022can).
Most prevalent poisoning attacks (NEURIPS2018_22722a34; EM; REM; Fang2019LocalMP) share a common limitation: their malicious intentions are highly apparent. The victim may notice a significant disparity in the performance of the model between the training and validation datasets. In response, strategies such as adversarial training (bai2021recent), appropriate data preprocessing methods (apbench), or the elimination of anomalous (liao2018defense) data can be employed as defensive measures against these attacks.
@@ Introduction
In this paper, we reveal a more threatening method of poisoning attack namely Deferred Poisoning Attack (DPA). As the term ”Deferred” implies, this attack does not disrupt the training process, allowing the model to maintain normal performance over the validation set. Instead, the ”toxicity” of the attack manifests by undermining the model’s robustness at the deployment stage. Fig. 1 illustrates the scenario of our attack by the case of Machine Learning as a Service (MLaaS) (ribeiro2015mlaas). Attackers can use DPA to train a model that performs normally during training and testing, correctly identifying signs like ”Speed Limit 50” to gain user trust. However, once deployed in systems like autonomous driving, this vulnerable model becomes susceptible to evasion attacks or even natural disturbances such as fog, rain, and lighting variations. This can cause critical failures, such as misclassifying a speed limit sign as ”STOP”.
We fulfill the purpose of DPA by forcing the model trained over the contaminated dataset to converge to a similar point as the one trained over the clean dataset, making the poisoned model perform normally on the validation dataset. On the other hand, we enlarge the local curvature of the poisoned model around each sample in the training dataset to amplify the sensitivity of the poisoned model. Fig. 2 illustrates the above motivation. A large local curvature (the red curve) results in a significant increase in model loss with a small perturbation of a given sample. In contrast, a small local curvature (the blue curve) enables the model loss to remain stable even with a large perturbation. Formally, a large local curvature implies that the Hessian matrix is ill-conditioned with a large conditional number (boyd2004convex). Along this line, our DPA generates poisoned samples to induce the model trained on this contaminated dataset to become singularization (a large conditional number.) with respect to the input samples.
@@
  • We validate the generality of the DPA across a broader range of scenarios.
@@
Related work
Poisoning attacks aim to insert malicious samples into training datasets, thereby compromising the ability of the resulting models (biggio2012poisoning; Biggio2017WildPT; Jagielski2018ManipulatingML; Liu2018TrojaningAO). EM first proposed a method for calculating perturbations by minimizing empirical risk, which impedes the models’ ability to extract valuable information from images. REM made a more robust improvement by minimizing adversarial training loss to obtain a more robust set of unlearnable examples. However, previous poisoning attacks directly compromised the integrity of the models, making it easy for us to detect anomalies in the model’s performance on the validation set during the training phase, which then allows us to take defensive measures. In this paper, we identify a form of deferred poisoning that appears normal during the model training phase yet renders the poisoned model extremely vulnerable. This kind of stealthiness can mislead researchers, thereby posing a serious threat to their work.
Deferred poisoning attack
In this section, we introduce the principles of DPA, detailing how to generate such perturbations and providing both the necessary theoretical and empirical analysis.
Problem statement
@@
We assume that adversaries have access to the entire training dataset in the MLaaS scenario but are unaware of the target model’s parameter structure or any output, and do not interfere with the target model’s training process.
@@
$\text{s.t.}\;{\|\delta\|}_{\infty}\leq\epsilon.$
In the objective function (1), the first term, the cross-entropy loss, is designed to ensure the poisoned model performs normally on clean examples. The second and third terms are dedicated to updating the poisoning perturbation $\delta$. Notably, our approach differs significantly from existing poisoning attack methods (wu2022one; tian2022comprehensive), which typically aim to increase the model loss on poisoned examples so as to degrade the model’s performance on the validation dataset. Our proposed loss still allows the model to converge on the poisoned examples but introduces a unique design: the Hessian Singularization term $Q(f_{\theta}(\hat{x}),y)=\textit{tr}(H^{T}H)$, where $H$ is the Hessian matrix of the loss function with respect to the input $\hat{x}$. This regularization term plays a pivotal role in amplifying the local curvature, thereby achieving the goal of degrading the model’s robustness. The value of $Q(f_{\theta}(\hat{x}),y)$ is positively related to the curvature around the sample $\hat{x}$. More of the relevant theoretical derivation about the Hessian Singularization term will be discussed in Session Hessian singularization. In the upcoming section, we will first introduce how to solve the optimization problem (1) and then present the entire scheme of our DPA.
Generating perturbation
Empirically, to ensure that the optimization of the problem (1) converges effectively, we divide this problem into two alternating iterative sub-optimization problems, which have been widely adopted in the research area of poisoning attack (munoz2017towards; REM). Specifically, as shown in Fig. 3, the first phase is the model training stage. We consider the following sub-problem
$\theta\leftarrow\mathop{\arg\min}_{\theta}\bigg[\mathcal{L}(f_{\theta}({x}),y)+\mathcal{L}(f_{\theta}(\hat{x}),y)\bigg].$ (2)
@@
Hessian singularization
In this section, we will explain why the Hessian Singularization could impact the robustness of deep models. Consider the loss function of a model denoted by $l$ and a given input $x$ and assume that this function is strictly convex around a small neighbor of the $x$. For any small perturbation $h$ around $x$, the loss function satisfies the following conditions (boyd2004convex):
@@
where $tr()$ represents the trace of a matrix (see (lancaster1985theory) for the details of this inequality). Consequently, this leads to the proposed Hessian Singularization regularization term $Q(f_{\theta}^{*}(x+\delta),y)=\text{tr}(H^{T}H)$ in the objective function (1).
@@
Relaxation of Hessian singularization term
@@
The new strategy avoids directly computing the Hessian matrix and significantly reduces the computational cost to approximately 0.008 seconds and still maintains the effectiveness of our DPA.
@@
Table 1: The comparison of the accuracy and robustness(%) of the clean model and poisoned model against different kinds of evasion attacks.
@@
Table 2: The performance of various poisoning attacks at different attack scales.
Experiments
Experiment setting
Datasets & Models
We generate perturbation on the CIFAR10 (krizhevsky2009learning), SVHN (netzer2011reading), and TinyImageNet (imagenet) datasets respectively. Subsequently, we use the VGG16 (simonyan2014very), ResNet18 (resnet), and ResNet50 (resnet) models to train on the CIFAR10 and SVHN datasets. For the TinyImageNet, we use the ResNet18, ResNet50, and DenseNet121 (huang2017densely) models for training. It should be noted that the TinyImageNet is a subset of the ImageNet dataset.
Evaluation metric
The major claim of our work is that models trained over a dataset poisoned by our DPA would exhibit a lower robustness against adversarial perturbation or even natural noises. Thus we quantize the robustness of a victim model by adopting the measure defined by (moosavi2016deepfool). That is,
where $\hat{r}(x)$ is the minimal perturbation for a successful attack, and $\mathscr{D}$ denotes the validation set. $p$ is $2$ (e.g., DeepFool (moosavi2016deepfool)) or $\infty$ (e.g., PGD). A smaller $\hat{\rho}(f)$ indicates a model is more vulnerable.
On the other hand, we emphasize that our DPA could significantly decrease the degradation of the model performance on clean examples, exhibiting the stealthiness of our DPA. To this end, we apply traditional poisoning methods such as EM, REM, LSP (LSP), and AR (AR), as well as our proposed method to contaminate datasets (apbench). We assess the classification accuracy of the poisoned models, the closer the accuracy is to that of the clean models, the more effective the stealth of the poisoning attack method.
Figure 5: Comparison of examples generated by different Poisoning attacks above the dotted line(i.e. EM, REM and LSP). For each attack, we show the poisoned sample (top) and the magnified (×5) residual (bottom).
Effectiveness of deferred poisoning attack
As shown in Table 1 where $\hat{\rho}_{F}$, $\hat{\rho}_{P}$ and $\hat{\rho}_{D}$ represent the robustness of victim models attacked by FGSM (FGSM), PGD (PGD), and DeepFool respectively, we can observe that the ACC of both the clean models and the poisoned models are almost the same. For example, the 6th and 7th rows, corresponding to the 2nd to 5th columns, illustrate the performance of the model trained with ResNet18 on TinyImageNet. After being subjected to DPA, the poisoned model maintains the same accuracy as the clean model on the validation set, both at 0.72. However, the robustness of the poisoned model against FGSM, PGD, and DeepFool adversarial attacks has significantly decreased from 0.74, 0.50, and 0.28 to 0.48, 0.42, and 0.09, respectively. Similarly, by examining the data for other models and datasets in the table, we arrive at the same conclusion.
Model Clean Gaussian Poisson Speckle Rayleigh
VGG16 C 0.89 0.80 0.71 0.72 0.82
P 0.75 0.45 0.47 0.48 0.44
ResNet18 C 0.81 0.60 0.63 0.62 0.66
P 0.76 0.41 0.49 0.41 0.43
ResNet50 C 0.78 0.63 0.61 0.65 0.68
P 0.73 0.37 0.50 0.38 0.42
Table 3: The impact of natural noise on the poisoned (P) and clean (C) models.
Transferability
It is important to note that, in these examples from Table 1, perturbations are generated based on the VGG16 model that was pretrained on the full ImageNet dataset. However, the toxic effects of these perturbations are not only effective on the VGG16 model (white-box scenario) but also transfer to other models such as ResNet18, ResNet50, and DenseNet121 (black-box scenarios). This indicates that DPA possesses excellent transferability, thereby making it highly suitable for black-box attack scenarios. The poisoned model being simultaneously affected by adversarial samples and natural noise indicates that DPA fundamentally reduces the model’s robustness to arbitrary perturbations. This corroborates our claim that by increasing the curvature of the loss function around local optima, the model becomes more vulnerable.
Generalized scenario
Adversarial examples do not frequently manifest in real-world scenarios. According to our claim, DPA fundamentally reduces the model’s robustness to arbitrary perturbations, that’s means poisoned model not only simultaneously affected by adversarial samples but also natural noise. To substantiate this, we introduce random noise sampled from different distributions to the test images to simulate more general conditions. We experimented with natural noise from various distributions: Gaussian Noise, Poisson Noise, Speckle Noise and Rayleigh Noise.
Table 3 compares the classification accuracy of clean and poisoned models on CIFAR10 with different sampled random noise injections. It is evident that the poisoned models are significantly more sensitive to random noise, indicating that our method can render the models vulnerable. For example, when perturbing VGG16 with Gaussian noise, the clean model achieves a high classification accuracy of 0.80, remaining virtually unaffected by the noise. However, when the same Gaussian noise is used to disturb the poisoned model, the classification accuracy drops to only 0.45. This indicates that DPA renders the model considerably vulnerable even to natural noise, and poses a significant potential threat to the field of artificial intelligence security.
Stealthiness of poisoned samples
In this section, we verify the low attack cost characteristic of DPA and thereby achieve the stealthiness of the poisoned samples. DPA has a significant advantage over traditional data poisoning methods, it requires only a minimal perturbation cost, such as $\epsilon=3/255$, to achieve excellent results. In contrast, data poisoning methods with $l_{\infty}$ norm constraints, such as EM and REM, do not demonstrate effectiveness at this level of perturbation.
By contrasting the classification accuracy of adversarial samples between the poisoned and clean models, we present the effectiveness of DPA in a more intuitive form. Table 2 presents the classification accuracy of the model against different adversarial samples after being subjected to poisoning attacks with varying perturbation radii, where all adversarial samples’ perturbation radii are set to 1/255. It shows that the EM and REM methods do not fully take effect when $\epsilon$ are as low as $3/255$ or even $5/255$ across different models, at least $\epsilon=8/255$ to exhibit aggressiveness. In contrast, our method shows excellent offensiveness even at the scale of $\epsilon=3/255$. For instance, in rows 3 to 6 of Table 2 for the VGG16 model trained on CIFAR10, when $\epsilon=3/255$, the models trained with EM and REM maintain high classification accuracy and accuracy post-adversarial attack. However, the model poisoned by DPA exhibits extremely low accuracy post-adversarial attack, indicating that the models are very vulnerable and can be significantly impacted by adversarial examples due to the approach mentioned. This also suggests that the attack cost of DPA is much lower than that of traditional poisoning attacks, while also ensuring the stealthiness of our method.
As depicted in Fig. 7, the perturbation we propose is difficult to discern with the naked eye. Compared to other poisoning attacks, even when the residual images between the poisoned and clean images are magnified five times, the noise is hardly noticeable. This illustrates the profound visual stealthiness inherent in our method, which is alarming in terms of its potential for misuse.
Comparison with backdoor attacks
@@
Models attacked by backdoor attacks behave normally during the training phase, but make mistakes when encounter trigges. However, DPA does not implant specific Trojans, more generally, its goal is to make poisoned models particularly vulnerable.
Methods ACC ASR $\hat{\rho}_{F}\downarrow$ $\hat{\rho}_{P}\downarrow$ $\hat{\rho}_{D}\downarrow$
@@
Ours 0.75 - 0.58 0.50 0.32
Table 4: The accuracy and robustness(%) of VGG16 after training CIFAR10 with backdoors implanted.
@@
Table 4 demonstrates that even when the attack success rate (ASR) is very high, backdoor attacks do not render the model more vulnerable. This indicates that the aggressiveness of DPA is insidious and unique, it can make the model more vulnerable while ensuring normal model performance, bringing unforeseen potential harm to machine learning.
Robustness analysis
@@
We conduct various data augmentation techniques (cutout; image_process) on data-infused with deferred poison. By observing the performance of models trained on these augmented datasets, we assess the robustness of DPA.
@@
Table 8 displays the performance of DPA when facing various data augmentation techniques. Standard refers to the basic DPA with no modifications. This table evident that our method can still render models vulnerable under most data augmentation methods, indicating that our approach possesses strong robustness.
Method ACC $\hat{\rho}_{F}\downarrow$ $\hat{\rho}_{P}\downarrow$ $\hat{\rho}_{D}\downarrow$
Clean 0.81 1.62 1.01 1.24
Standard 0.76 0.79 0.57 0.39
Pretrained 0.75 1.12 0.59 0.42
Cutout 0.77 0.92 0.60 0.44
Gaussian 0.78 0.87 0.58 0.42
Gray 0.73 0.62 0.53 0.35
JPEG 0.79 1.16 0.64 0.60